# Clarifying identification and estimation of treatment effects in the Sequential Parallel Comparison Design

### v1

@@ 5 Discussion
In summary, we derived the estimands for each subset of participants (all participants, placebo non-responders, and placebo responders) in the SPCD while characterizing the placebo non-responder status as a latent binary variable. The latent status is estimated by classifier $C_{R}(y_{0},y_{1})$ using the baseline and stage 1 data for the placebo arm. We demonstrate that two of the conventional estimators for the SPCD, $\hat{\theta}_{2}$ (Equation 11) and $\hat{\theta}_{w},$ (Equation 13) do not target their intended estimands while $\hat{\theta}_{1}$ does unbiasedly estimate $\Delta_{\mathrm{all}}.$ Our analytical and simulation results indicate that placebo non-response misclassification is a central cause of the bias.
@@
From Equation 12, the conditions for which $E(\hat{\theta}_{2})=\Delta_{NR}$, the desired expectation, are (i) $P(L=1|R=0)=0,$ equivalently, $R_{i}=0\Rightarrow L_{i}=0$ for all placebo non-responders (i.e. the classifier only identifies true non-responders as non-responders), or (ii) $\Delta_{\mathrm{placebo}}=0$. However, both of these conditions are unverifiable. In the first condition (i) the probability of placebo responder misclassification, $P(L=1|R=0)$, depends on the unknown latent status $L,$ so we cannot check the negative predictive performance of $C_{R}(y_{1},y_{0})$. The second condition (ii) requires that the placebo has no effect. This condition is not testable, is not typically a reasonable assumption, and defeats the purpose of using an SPCD to estimate the placebo effect. Condition (i), the assumption of perfect classification, is typically assumed implicitly in order to obtain an unbiased estimator without acknowledging that it is impossible to check if it is even plausible[rybin2018, chi2015, cui2019, rybin2015].
@@
The weighted estimator, $\hat{\theta}_{w}$, has been the primary inferential tool in the SPCD literature[chen2011, rybin2018, tamura2011, chi2015, rybin2015], despite the conceptual difficulty in interpreting its expectation given by Equation 14. Primarily, it is unclear how to interpret the weighted average component (wP(L=1)+(1w)P(L=1|R=0))(wP(L=1)+(1-w)P(L=1|R=0)) of the expectation of $\hat{\theta}_{w}$. This conceptual challenge can be resolved under the above implausible conditions for unbiasedness for $\hat{\theta}_{2}$ or under the third condition (iii) that $P(L=1|R=0)=P(L=1).$
@@
Under condition (iii), the expectation simplifies to $E(\hat{\theta}_{w})=\Delta_{NR}-P(L=1)\Delta_{\mathrm{placebo}}=\Delta_{\mathrm{all}}.$ However, $P(L=1|R=0)=P(L=1)$ occurs if $R\perp L$, $P(L=1)=0$, $\Delta_{\mathrm{placebo}}=0$, or $w=1$ (see Result 9). All four conditions defeat the purpose of using the SPCD. The classified placebo responder status $R$ being independent of $L$ implies that we do not gain any information by performing the classification and stage 2 is essentially a replication of stage 1 because true placebo responders are not excluded from stage 2. The second and third conditions $(P(L=1)=0\text{ or }\Delta_{\mathrm{placebo}})$ imply that there are no placebo responders in the population to begin with. Lastly, choosing $w=1$ results in only using the stage 1 data which reduces the SPCD to a simple randomized controlled trial.
@@
If we are trying to estimate $\Delta_{NR},$ then we should use $w=0\Rightarrow\hat{\theta}_{w}=\hat{\theta}_{2}$ and hope that $P(L=1|R=0)$ is small (see Result 10). When $w=0,$ the estimation only uses placebo arm stage 1 data to determine $C_{R}(y_{0},y_{1})$ and throws away the entire active treatment arm where $A_{1}=1.$ In order to accurately estimate $\Delta_{NR}$ with $\hat{\theta}_{2}$, whichever (deterministic or stochastic) classifier $C_{R}(y_{0},y_{1})$ that is used to determine responder status $R$ must perfectly identify true responders $L=1.$ If that is not the case then some true responders $L=1$ may be classified as non-responders by $R=0$ creating a mixture at Stage 2.
@@
Because the true latent statuses are unknown, there is uncertainty in the classifications. In practice, we should require the inferential model or estimators to account for the classification uncertainty in the variance estimate of the stage 2 estimator. However, the classification uncertainty is unaccounted for in previous definitions of the SPCD estimator[chi2015, cui2019, liu2022]. The previously shown suboptimal control of Type I error by the weighted SPCD estimator $\hat{\theta}_{w}$ may be partially explained by the stage 2 estimator failing to unbiasedly estimate $\Delta_{NR}$ due to the misclassification[cui2019, liu2022].
@@
The overall weighted estimator $\hat{\theta}_{w}$ is particularly problematic. Using any $w\neq 0,1$ results in a biased weighted estimator $\hat{\theta}_{w}$ whose expectation (Equation 14) does not correspond to a quantity of clear clinical interest. The placebo non-responder ATE is adjusted for placebo response $\Delta_{\mathrm{placebo}}$ by some weighted average of the proportion of placebo responders in the population and the proportion of responders that are misclassified as non-responders. It does not seem likely that this quantity would be of clinical interest or importance given that it is a function of the statistical analysis through the classifier $C_{R}(y_{0},y_{1}).$
@@
Our results are related to prior work[liu2022], but highlight the importance of the accuracy of classification and provide a quantification of the bias of estimates due to misclassification. Where the previous work[liu2022] uses a counterfactual variable to define the responder status so that both drug and placebo responders and non-responders are defined resulting in four strata, we define only placebo response through our non-counterfactual, latent status $L$ which assumes that everyone would be a drug responder. This in effect collapses their never-responder and drug-only responder strata[liu2022], while the always-responder strata correspond to our placebo responders and active treatment participants.
Additionally, Liu et al.[liu2022] focus on the estimation of the individual strata effects at each time point, and the effects for never-responders and drug-only responders by taking a weighted average of their respective estimates at each stage. Their results assume that classification is accurate following the EM algorithm’s maximum likelihood estimation and sampling of responder status. Our results provide insight of the impact of misclassification on the estimates.
Beyond the scope of this investigation, it may be better to model placebo response as a continuous latent variable rather than a binary indicator since it is plausible that everyone can exhibit some degree of placebo response along a continuum. While the binary responder status is a simplifying assumption that ties estimation to the well-known finite mixture model, it may not be a reasonable assumption for many of the applications in which SPCD are deployed, e.g., to study interventions for pain or psychiatric conditions.
@@
Estimand Derivations
We provide identification results for each of the stage-1 estimands under the structural causal model described in Section 2, assuming Assumptions 16.
Result 1 (Average Placebo Effect).
The average effect due to placebo is
$\Delta_{\text{placebo}}=E(Y_{1}^{a_{1}=0,l=1})-E(Y_{1}^{a_{1}=0,l=0}).$
Proof.
First, we establish the identification of the expectation of the counterfactuals from the observed data. For $l\in\{0,1\},$
$E(Y_{1}^{a_{1}=0,l})$ $=E(Y_{1}^{a_{1}=0,l}\mid A_{1}=0)$ by Assumption 2
$=E(Y_{1}^{a_{1}=0,l}\mid A_{1}=0,L=l)$ by Assumption 4
$=E(Y_{1}\mid A_{1}=0,L=l)$ by Assumptions 1, 3
$\text{identified from }f_{Y_{1}\mid A_{1}=0}(y)$ by Assumption 6
Then,
$E(Y_{1}^{a_{1}=0,l=1}$ $-Y_{0})-E(Y_{1}^{a_{1}=0,l=0}-Y_{0})=\left[E(Y_{1}^{a_{1}=0,l=1})-E(Y_{0})\right]-\left[E(Y_{1}^{a_{1}=0,l=0})-E(Y_{0})\right]$
$=E(Y_{1}^{a_{1}=0,l=1})-E(Y_{1}^{a_{1}=0,l=0})$
$=\Delta_{\mathrm{placebo}}.$ (16)
Note that $E(Y^{a_{1}=0,l=1})=\Delta_{\mathrm{placebo}}+E(Y^{a_{1}=0,l=0}).$
Result 2 (Average Treatment Effect for Placebo Non-Responders).
The average treatment effect for true placebo non-responders is
$\Delta_{\text{NR}}=E[Y_{1}^{a_{1}=1,l=0}]-E[Y_{1}^{a_{1}=0,l=0}].$
Proof.
Again, we begin by identifying the counterfactual’s expectations from the data.
$E(Y_{1}^{1,0})$ $=E(Y_{1}^{1})$ by Assumption 5
$=E(Y_{1}\mid A_{1}=1)$ by Assumptions 1, 2, 3
$E(Y_{1}^{0,0})$ $=E(Y_{1}\mid A_{1}=0,L=0)$ by Assumptions 1, 2, 3, 4
$\text{identified from }f_{Y_{1}\mid A_{1}=0}(y)$ by Assumption 6
Then,
$E(Y_{1}^{a_{1}=1}-Y_{0})$ $-E(Y_{1}^{a_{1}=0,l=0}-Y_{0})=\left[E(Y_{1}^{a_{1}=1})-E(Y_{0})\right]-\left[E(Y_{1}^{a_{1}=0,l=0})-E(Y_{0})\right]$
$=E(Y_{1}^{a_{1}=1})-E(Y_{1}^{a_{1}=0,l=0})$
$=\Delta_{NR}.$ (17)
Note that $E(Y_{1}^{a_{1}=1})-E(Y_{1}^{a_{1}=0,l=0})=\Delta_{NR}.$
Result 3 (Average Treatment Effect for Placebo Responders).
The average treatment effect for true placebo responders is
$\Delta_{\text{PR}}=E(Y_{1}^{a_{1}=1,l=1})-E(Y_{1}^{a_{1}=0,l=1}).$
We can also summarize the ATE for placebo responders as the quantity $\Delta_{PR}=\Delta_{NR}-\Delta_{placebo}.$
Proof.
$E(Y_{1}^{a_{1}=1,l=1})$ $=E(Y_{1}^{a_{1}=1,l=0})=E(Y_{1}\mid A_{1}=1)$ by Assumptions 5, 1, 2, 3
$E(Y_{1}^{a_{1}=0,l=1})$ $=E(Y_{1}\mid A_{1}=0,L=1)$ by Assumptions 1, 2, 3, 4
$\text{identified from }f_{Y_{1}\mid A_{1}=0}(y)$ by Assumption 6
$\Delta_{PR}$ $=E(Y_{1}^{a_{1}=1})-E(Y_{1}^{a_{1}=0,l=1})$
$=E(Y_{1}^{a_{1}=1})-\left[E(Y_{1}^{a_{1}=0,l=0})+\Delta_{\mathrm{placebo}}\right]$
[…]
Result 4 (Average Treatment Effect for All Participants - Difference).
The average treatment effect for all participants can be represented as a simple difference
$\Delta_{\text{all}}=E(Y_{1}^{a_{1}=1})-E(Y_{1}^{a_{1}=0}).$
Proof.
$E(Y_{1}^{a})$ $=E(Y_{1}\mid A_{1}=a)$ by Assumptions 1, 2, 3
$\Rightarrow\Delta_{\text{all}}$ $=E(Y_{1}\mid A_{1}=1)-E(Y_{1}\mid A_{1}=0).$
Result 5 (Average Treatment Effect for All Participants - Mixture).
The ATE for all participants can also be represented as a mixture
$\Delta_{\text{all}}\;=\;(1-P(L=1))\Delta_{\text{NR}}+P(L=1)\Delta_{\text{R}}\;=\;\Delta_{\text{NR}}-P(L=1)\Delta_{\text{placebo}}.$
Proof.
Next, the ATE for the full population is presented in mixture form. $P(L=1)$ is identified from $f_{Y_{1}\mid A_{1}=0}(y)$ by Assumptions 3 and 6. $\Delta_{\text{NR}},\;\Delta_{\text{PR}},\;\Delta_{\text{placebo}}$ are identified as in Results 1-3. Then from algebra,
$E(Y_{1}^{a_{1}=1})-E(Y_{1}^{a_{1}=0,l})$ $=E(Y_{1}^{a_{1}=1})-E\left(P(L=1)Y_{1}^{a_{1}=0,l=1}+(1-P(L=1))Y_{1}^{a_{1}=0,l=0}\right)$
$=E(Y_{1}^{a_{1}=1})-\left[P(L=1)E(Y_{1}^{a_{1}=0,l=1})+(1-P(L=1))E(Y_{1}^{a_{1}=0,l=0})\right]$
$=E(Y_{1}^{a_{1}=1})-P(L=1)(\Delta_{\mathrm{placebo}}+E(Y_{1}^{a_{1}=0,l=0}))-(1-P(L=1))E(Y_{1}^{a_{1}=0,l=0})$
$=E(Y_{1}^{a_{1}=1})-E(Y_{1}^{a_{1}=0,l=0})-P(L=1)\Delta_{\mathrm{placebo}}$
$=\Delta_{NR}-P(L=1)\Delta_{\mathrm{placebo}}.$ (19)
Estimator Derivations
Result 6 (Expectation of Stage 1 Estimator).
$E(\hat{\theta}_{1})=\Delta_{\mathrm{all}}.$
Proof.
We make Assumptions 1-7. Then,
$E(\hat{\theta}_{1})$ $=E\left[\frac{1}{n_{A}}\sum_{i=1}^{n}(Y_{i,1}-Y_{i,0})I(A_{i,1}=1)-\frac{1}{n_{P}}\sum_{i=1}^{n}(Y_{i,1}-Y_{i,0})I(A_{i,1}=0)\right]$
$=\frac{1}{n_{A}}\sum_{i=1}^{n}E(Y_{i,1}-Y_{i,0}|A_{i,1}=1)-\frac{1}{n_{P}}\sum_{i=1}^{n}E(Y_{i,1}-Y_{i,0}|A_{i,1}=0)$
$=\frac{1}{n_{A}}\sum_{i=1}^{n_{A}}E(Y_{i,1}^{a_{1}=1}-Y_{i,0})-\frac{1}{n_{P}}\sum_{i=1}^{n_{P}}E(Y_{i,1}^{a_{1}=0,l}-Y_{i,0})$
$=E(Y_{1}^{a_{1}=1}-Y_{0})-E(Y_{1}^{a_{1}=0,l}-Y_{0})$
$=E(Y_{1}^{a_{1}=1})-E(Y_{1}^{a_{1}=0,l})$
$=\Delta_{\mathrm{all}}.$ (20)
Result 7 (Expectation of the Stage 2 Estimator).
$E(\hat{\theta}_{2})=\Delta_{NR}-q_{1}\Delta_{\mathrm{placebo}}.$
$E(\hat{\theta}_{2})=\Delta_{NR}$ if $q_{1}=P(L=1|R=0)=0,$ i.e. if no true placebo responders $L=1$ are classified as placebo non-responders $R=0.$
Proof.
Let $P(L=0,R=0)=p_{00}$ and $P(L=1,R=0)=p_{10}$ so that $q_{l}=P(L=l|R=0)=p_{l0}/(p_{00}+p_{10}).$ To calculate these probabilities, we would need to make a distributional assumption about $Y_{t}.$
We also define
$E(Y_{2}^{a_{1},a_{2},l}|R=0)$ $=P(L=0|R=0)E(Y_{2}^{a_{1},a_{2},l=0}|R=0)$
$~~~~~~~~+P(L=1|R=0)E(Y_{2}^{a_{1},a_{2},l=1}|R=0)$
$=q_{0}E(Y_{2}^{a_{1},a_{2},l=0}|R=0)+q_{1}E(Y_{2}^{a_{1},a_{2},l=1}|R=0)$ (21)
Again we make Assumptions 1-7 (modified accordingly for Stage 2), so that the placebo non-responder average treatment effect is the same as from Stage 1, i.e.
$E(Y_{2}^{a_{1}=0,a_{2}=1,l=0})-E(Y_{2}^{a_{1}=0,a_{2}=0,l=0})=E(Y_{1}^{a_{1}=1})-E(Y_{1}^{a_{0}=0,l=0})=\Delta_{NR}.$
Note that given $R$, the sample sizes for each stage 2 arm are fixed $n_{PA}$ and $n_{PP}$ since the classified responder status determines the samples size of each arm. Then we get
$E(\hat{\theta}_{2})=$ $E\left[\frac{1}{n_{PA}}\sum_{i=1}^{n}(Y_{i,2}-Y_{i,1})I(A_{i,1}=0,R_{i}=0,A_{i,2}=1)|R\right]$
$~~~~-E\left[\frac{1}{n_{PP}}\sum_{i=1}^{n}(Y_{i,2}-Y_{i,1})I(A_{i,1}=0,R_{i}=0,A_{i,2}=0)|R\right]$
$=\frac{1}{n_{PA}}\sum_{i=1}^{n_{PA}}E\left[Y_{i,2}-Y_{i,1}|A_{i,1}=0,R_{i}=0,A_{i,2}=1,\right]$
$~~~~-\frac{1}{n_{PP}}\sum_{i=1}^{n_{PP}}E\left[Y_{i,2}-Y_{i,1}|A_{i,1}=0,R_{i}=0,A_{i,2}=0\right]$
$=\frac{n_{PA}}{n_{PA}}E\left[Y_{2}^{a_{1}=0,a_{2}=1,l}-Y_{1}^{a_{1}=0,l}|R=0\right]$
$~~~~-\frac{n_{PP}}{n_{PP}}E\left[Y_{2}^{a_{1}=0,a_{2}=0,l}-Y_{1}^{a_{1}=0,l}|R=0\right]$
$=E\left[Y_{2}^{a_{1}=0,a_{2}=1,l}-Y_{1}^{a_{1}=0,l}|R=0\right]$
$~~~~-E\left[Y_{2}^{a_{1}=0,a_{2}=0,l}-Y_{1}^{a_{1}=0,l}|R=0\right]$
$=E\left[Y_{2}^{a_{1}=0,a_{2}=1,l}|R=0\right]-E\left[Y_{2}^{a_{1}=0,a_{2}=0,l}|R=0\right]$
$=E\left[Y_{2}^{a_{1}=0,a_{2}=1,l}|R=0\right]-E\left[Y_{2}^{a_{1}=0,a_{2}=0,l}|R=0\right]$
$=\left[q_{0}E(Y_{2}^{a_{1}=0,a_{2}=1,l=0}|R=0)+q_{1}E(Y_{2}^{a_{1}=0,a_{2}=1,l=1}|R=0)\right]$
$~~~~-\left[q_{0}E(Y_{2}^{a_{1}=0,a_{2}=0,l=0}|R=0)+q_{1}E(Y_{2}^{a_{1}=0,a_{2}=0,l=1}|R=0)\right]$
$=q_{0}\left[E(Y_{2}^{a_{1}=0,a_{2}=1,l=0}|R=0)-E(Y_{2}^{a_{1}=0,a_{2}=0,l=0}|R=0)\right]$
$~~~~+q_{1}\left[E(Y_{2}^{a_{1}=0,a_{2}=1,l=1}|R=0)-E(Y_{2}^{a_{1}=0,a_{2}=0,l=1}|R=0)\right]$
$=q_{0}\Delta_{NR}+q_{1}\Delta_{PR},~~~~\mbox{where }\Delta_{PR}=\Delta_{NR}-\Delta_{\mathrm{placebo}}\mbox{ is the placebo response.}$
$=q_{0}\Delta_{NR}+q_{1}(\Delta_{NR}-\Delta_{\mathrm{placebo}})$
$=\Delta_{NR}(q_{0}+q_{1})-q_{1}\Delta_{\mathrm{placebo}}$
$\Rightarrow E(\hat{\theta}_{2})$ $=\Delta_{NR}-q_{1}\Delta_{\mathrm{placebo}}\neq\Delta_{\mathrm{all}}.$ (22)
Result 8 (Expectation of the Weighted SPCD Estimator).
$E(\hat{\theta}_{w})=\Delta_{NR}-\Delta_{\mathrm{placebo}}(wP(L=1)+(1-w)P(L=1|R=0)).$
Proof.
$E(\hat{\theta}_{w})$ $=wE(\hat{\theta}_{1,\mathrm{all}})+(1-w)E(\hat{\theta}_{2})$
$=w\Delta_{\mathrm{all}}+(1-w)\left[\Delta_{NR}-q_{1}\Delta_{\mathrm{placebo}}\right]$
$=w(\Delta_{NR}-p_{L}\Delta_{\mathrm{placebo}})+(1-w)(\Delta_{NR}-q_{1}\Delta_{\mathrm{placebo}})$
$=\Delta_{NR}(w+(1-w))-\Delta_{\mathrm{placebo}}(wp_{L}+(1-w)q_{1})$
$=\Delta_{NR}-\Delta_{\mathrm{placebo}}(wP(L=1)+(1-w)P(L=1|R=0)).$ (23)
Result 9 (Weighted SPCD Estimator Targets $\Delta_{\mathrm{all}}$).
$\hat{\theta}_{w}$ is unbiased for $\Delta_{\mathrm{all}}$ if any of the following hold:
  1. 1.
    $P(L=1)=0.$
  2. 2.
    $\Delta_{\mathrm{placebo}}=0.$
  3. 3.
    $R\perp\!\!\!\perp L.$
  4. 4.
    We set $w=1.$
Proof.
  1. [font=]
  2. 1.
    Suppose $P(L=1)=0$.
    By Result 5, $\Delta_{\mathrm{all}}=\Delta_{NR}.$ Then,
    $E(\hat{\theta}_{w})$ $=\Delta_{NR}-\Delta_{\mathrm{placebo}}(wP(L=1)+(1-w)P(L=1|R=0))$
    $=\Delta_{NR}-\Delta_{\mathrm{placebo}}(w0+(1-w)0)~~~~\mbox{since }P(L=1)=0\Rightarrow P(L=1|R=0)=0$
    $=\Delta_{NR}=\Delta_{\mathrm{all}}.$ (24)
  3. 2.
    Suppose $\Delta_{\mathrm{placebo}}=0$.
    By Result 5, $\Delta_{\mathrm{all}}=\Delta_{NR}$. Then,
    $E(\hat{\theta}_{w})$ $=\Delta_{NR}-0(wP(L=1)+(1-w)P(L=1|R=0))$
    $=\Delta_{NR}=\Delta_{\mathrm{all}}.$ (25)
  4. 3.
    Suppose $R\perp\!\!\!\perp L$.
    Then $P(L=1|R=0)=P(L=1)=p_{L},$ and,
    $E(\hat{\theta}_{w})$ $=\Delta_{NR}-\Delta_{\mathrm{placebo}}(wP(L=1)+(1-w)P(L=1|R=0))$
    $=\Delta_{NR}-\Delta_{\mathrm{placebo}}(wp_{L}+(1-w)p_{L})$
    $=\Delta_{NR}-p_{L}\Delta_{\mathrm{placebo}}$
    $=\Delta_{\mathrm{all}}.$ (26)
    Note $R\perp\!\!\!\perp L$ is trivially true if $C_{R}$ assigns every participant $R_{i}=0.$
  5. 4.
    Suppose we set $w=1$.
    Then $\hat{\theta}_{w}=\hat{\theta}_{1}.$ See Result 6 for the proof $\hat{\theta}_{1}$ is unbiased for $\Delta_{\mathrm{all}}.$
Result 10 (Weighted SPCD Estimator Targets $\Delta_{NR}$).
$\hat{\theta}_{w}$ is unbiased for $\Delta_{NR}$ if $w=0\Rightarrow\hat{\theta}_{w}=\hat{\theta}_{2}$ and $\hat{\theta}_{2}$ is unbiased for $\Delta_{NR}.$
Proof.
See Result 7 for proof of when $\hat{\theta}_{2}$ is unbiased for $\Delta_{NR}.$

### v2

@@ 6 Discussion
In summary, we derived the estimands for each subset of participants (all participants, placebo non-responders, and placebo responders) in the SPCD while characterizing the placebo non-responder status as a latent binary variable. The latent status is estimated by the classifier $C_{R}(y_{0},y_{1})$ using the baseline and stage 1 data for the placebo arm. Further, we demonstrated that two of the conventional estimators for the SPCD, $\hat{\theta}_{2}$ (Equation 11) and $\hat{\theta}_{w},$ (Equation 13) do not target their intended estimands while $\hat{\theta}_{1}$ does unbiasedly estimate $\Delta_{\mathrm{all}}.$ Our analytical and simulation results indicated that placebo non-response misclassification is a central cause of the bias.
@@
Specifically from Equation 12, the conditions for which $E(\hat{\theta}_{2})=\Delta_{NR}$, the desired expectation, are (i) $P(L=1|R=0)=0,$ equivalently, $R_{i}=0\Rightarrow L_{i}=0$ for all placebo non-responders (i.e. the classifier only identifies true non-responders as non-responders), or (ii) $\Delta_{\mathrm{placebo}}=0$. However, both of these conditions are unverifiable. In the first condition (i) the probability of placebo responder misclassification, $P(L=1|R=0)$, depends on the unknown latent status $L,$ so we cannot check the negative predictive performance of $C_{R}(y_{1},y_{0})$. The second condition (ii) requires that the placebo has no effect. This condition is not testable, is not typically a reasonable assumption, and defeats the purpose of using an SPCD to estimate the placebo effect. Condition (i), the assumption of perfect classification, is typically assumed implicitly in order to obtain an unbiased estimator without acknowledging that it is impossible to check if it is even plausible[15, 2, 3, 14].
@@
Further, the weighted estimator, $\hat{\theta}_{w}$, has been the primary inferential tool in the SPCD literature[1, 15, 18, 2, 14], despite the conceptual difficulty in interpreting its expectation given by Equation 14. Primarily, it is unclear how to interpret the weighted average component (wP(L=1)+(1w)P(L=1|R=0))(wP(L=1)+(1-w)P(L=1|R=0)) of the expectation of $\hat{\theta}_{w}$. This conceptual challenge can be resolved under the above implausible conditions for unbiasedness for $\hat{\theta}_{2}$ or under the third condition (iii) that $P(L=1|R=0)=P(L=1).$
@@
Under condition (iii), the expectation simplifies to $E(\hat{\theta}_{w})=\Delta_{NR}-P(L=1)\Delta_{\mathrm{placebo}}=\Delta_{\mathrm{all}}.$ However, $P(L=1|R=0)=P(L=1)$ occurs if $R\perp L$, $P(L=1)=0$, $\Delta_{\mathrm{placebo}}=0$, or $w=1$ (see Result 9). All four conditions defeat the purpose of using the SPCD. The classified placebo responder status $R$ being independent of $L$ implies that we do not gain any information by performing the classification and stage 2 is essentially a replication of stage 1 because true placebo responders are not excluded from stage 2. The second and third conditions $(P(L=1)=0\text{ or }\Delta_{\mathrm{placebo}}=0)$ imply that no placebo responders are in the population to begin with. Lastly, choosing $w=1$ results in only using the stage 1 data which reduces the SPCD to a traditional randomized two-arm trial.
@@
If we are trying to estimate $\Delta_{NR},$ then we should use $w=0\Rightarrow\hat{\theta}_{w}=\hat{\theta}_{2}$ and hope that $P(L=1|R=0)$ is small (see Result 10). When $w=0,$ the estimation only uses placebo arm stage 1 data to determine $C_{R}(y_{0},y_{1})$, and throws away the entire active treatment arm where $A_{1}=1.$ In order to accurately estimate $\Delta_{NR}$ with $\hat{\theta}_{2}$, whichever (deterministic or stochastic) placebo response classifier $C_{R}(y_{0},y_{1})$ that is used must perfectly identify true responders $L=1.$ If that is not the case then some true responders $L=1$ may be classified as non-responders by $R=0$ creating a mixture at Stage 2.
@@
Because the true latent statuses are unknown, there is uncertainty in the classifications. In practice, we should require the inferential model or estimators to account for the classification uncertainty in the variance estimate of the stage 2 estimator. However, the classification uncertainty is unaccounted for in previous definitions of the SPCD estimator[2, 3, 9, 16]. The previously shown suboptimal control of Type I error by the weighted SPCD estimator $\hat{\theta}_{w}$ may be partially explained by the stage 2 estimator failing to unbiasedly estimate $\Delta_{NR}$ due to the misclassification[3, 9].
@@
Of key importance to recognize is that the overall weighted estimator $\hat{\theta}_{w}$ is particularly problematic. Using any $w\neq 0,1$ results in a biased weighted estimator $\hat{\theta}_{w}$ whose expectation (Equation 14) does not correspond to a quantity of clear clinical interest. The placebo non-responder ATE is adjusted for placebo response $\Delta_{\mathrm{placebo}}$ by some weighted average of the proportion of placebo responders in the population and the proportion of responders that are misclassified as non-responders. It does not seem likely that this quantity would be of clinical interest or importance given that it is a function of the statistical analysis through the classifier $C_{R}(y_{0},y_{1}).$
@@
Our results are related to prior work[9, 16], but highlight the importance of the accuracy of classification and provide a quantification of the bias of estimates due to misclassification. Where the previous work[9] uses a counterfactual variable to define the responder status so that both drug and placebo responders and non-responders are defined resulting in four strata, we define only placebo response through our non-counterfactual, latent status $L$ which assumes that everyone would be a drug responder, similar to the approach from Song et al.[16] This in effect collapses Liu et al.’s never-responder and drug-only responder strata[9], while the always-responder strata correspond to our placebo responders and active treatment participants.
Additionally, Liu et al.[9] focus on the estimation of the individual strata effects at each time point, and the effects for never-responders and drug-only responders by taking a weighted average of their respective estimates at each stage. Their results assume that classification is accurate following the EM algorithm’s maximum likelihood estimation and sampling of responder status. Our results provide insight of the impact of misclassification on the estimates.
While Song et al.[16] advocate for using their newly proposed estimands and estimators with a novel approach to classification, they acknowledge that misclassification will result in bias even for their estimators, and that misclassification is unavoidable. Additionally, despite acknowledging the difficulty in interpreting the SPCD estimator $\hat{\Delta}_{w}$ because it mixes study populations, Song et al. advocates for adopting the $\Delta_{Joint}$ estimand which also mixes study populations in an uninterpretable manner (participants in the first stage vs participants in the second stage).
Given the consensus on the importance of accurate placebo response classification in the inferential procedure for SPCD analysis, the intractability of placebo response misclassification, and the demonstrated amount of bias due to misclassification shown in Section Estimator Derivations as well as in prior work[9, 16], we would advocate for using alternative approaches to account for placebo response in clinical trials such as optimal partitioning via finite mixture models[19] or latent regression analysis where placebo response is reconsidered as a continuous latent predictor[20].
Beyond the scope of this investigation, it may be better to model placebo response as a continuous latent variable rather than a binary indicator since it is plausible that everyone can exhibit some degree of placebo response along a continuum. While the binary responder status is a simplifying assumption that ties estimation to the well-known finite mixture model, it may not be a reasonable assumption for many of the applications in which SPCD are deployed, such as in the study of interventions for pain or psychiatric conditions.